OpenStudy (anonymous):

-4(x + 10)-6x= -3(x - 2)-5

OpenStudy (anonymous):

Alright, so we need to remember PEDMAS

OpenStudy (anonymous):

And the Order of Operations

OpenStudy (anonymous):

Right? :)

OpenStudy (anonymous):

So, -4(x + 10)-6x= -3(x - 2)-5 , what we would do first is the Parenthesis

OpenStudy (anonymous):

-4(x+10)= -4x+-40

OpenStudy (anonymous):

and 3(x-2)= 3x -6

OpenStudy (anonymous):

So, now our equation looks like this: -4x -40 -6x = 3x-6 x 5

OpenStudy (anonymous):

im still confused

OpenStudy (anonymous):

Ok, hold on one sec, k?

OpenStudy (anonymous):

ok thanks

OpenStudy (anonymous):

Ok, so now we have -4x -40 -6x = 3x-6 x 5

OpenStudy (anonymous):

So, we need to get the variable isolated

OpenStudy (anonymous):

ok what does that mean lolol soory

OpenStudy (anonymous):

haha, it means we need to get X on one side and the numbers on the other

OpenStudy (anonymous):

We can combine some like terms here, so you see where? Hint: like terms are like 3x - 5x or 4+6 ;)

OpenStudy (anonymous):

okay so how would I do that? ;)

OpenStudy (anonymous):

Do you know how to add like terms?

OpenStudy (anonymous):

no im in 7th grd this is practice for hs

OpenStudy (anonymous):

Oh, ok :) Alright.

OpenStudy (anonymous):

-4x -40 -6x = 3x-6 x 5 OR we could write it as: -4x -6x -40 = 3x -6 x 5

OpenStudy (anonymous):

what is 3x - 6 x 5? It would be 3x -30 right?

OpenStudy (anonymous):

YEAH I somewhat get it

OpenStudy (anonymous):

Alright so on one side we have -4x -6x -40 and on the other we have 3x - 30

OpenStudy (anonymous):

@marichuiolopez do you want to do the rest, or do you want me to keep working thorugh it?

OpenStudy (anonymous):

can I get the rest please

OpenStudy (anonymous):

Alright, no problem :)

OpenStudy (anonymous):

thank you!

OpenStudy (anonymous):

So, -4x -6x -40. Well, lets combine like terms, so -4x -6x = -10x so now we have -10x -40

OpenStudy (anonymous):

Now our equation is: -10x - 40 = -3x -30

OpenStudy (anonymous):

We want to have a positive variable, so we are going to add -3x to -10x

OpenStudy (anonymous):

So, now we have: -7x -40 = -30

OpenStudy (anonymous):

lets add -40 to -30

OpenStudy (anonymous):

now we have -7x= 10

OpenStudy (anonymous):

we are going to divide both sides by -7, so the answer is 10/-7

OpenStudy (anonymous):



OpenStudy (anonymous):

Make sense? :)

OpenStudy (anonymous):

oh ok thanks for everything(:

OpenStudy (anonymous):

Haha, no problem :)